OpenStudy (anonymous):

catarina scored 89 and 84 on her last two tests. To raise her mean score to 90, what must she score on her next test?

OpenStudy (chriss):

\[\frac{89+84+x}{3}=90\] 173+x=270 x=270-173 x=97

OpenStudy (anonymous):

thanks

OpenStudy (chriss):

yw

OpenStudy (anonymous):

:)